# CryptOracle: A Modular Framework to Characterize FHE

### v2

@@ 4.1 Software Setup
To understand the trade-offs between computational efficiency and security, we sweep the ciphertext ring dimension N across and the computational depth L across [10, 20]. We also profile the predefined security standards (128, 192, and 256-bit security from [5]), which override the default parameter selection. To ensure a fair comparison, security parameters, evaluation keys, and input ciphertexts are generated once, serialized to files, and then reused by all primitives and microbenchmarks. Each workload fixes any additional application-specific parameters. Table 5 lists the default settings used in the evaluation, unless stated otherwise. During data collection, the median of five runs is reported for each metric. For primitive profiling, where an individual operation may complete in tens of microseconds, we invoke each primitive repeatedly until the cumulative runtime exceeds 500ms, then compute per-call averages based on the total number of invocations.
| Application | Sec. Std. | N | Batch Size | L |
| Primitives | none | 10 | ||
| Microbenchmarks | none | 10 | ||
| cifar10 | none | 5 | ||
| chi-square | none | 1 | 3 | |
| resnet | none | 10 | ||
| logreg | 128-bit | 14 |
@@ 5.1.1 Primitive Analysis
Figure 6 explores the microarchitectural behaviour of the primitives on the AMD platform, presenting a subset of the events listed in Table 4. At low thread counts, EvalMult and EvalRotate record the largest instruction footprints, underscoring their computational intensity. As the thread count rises from one to four, the instruction counts drop substantially, most notably for EvalMult, indicating that multithreading removes redundant serial work. In contrast, cache references increase with additional threads because of greater data sharing. The pattern of dTLB loads follows the cache trend, as increased parallel memory accesses raise TLB pressure and introduce page-level contention. Correspondingly, page-fault rates fall as more threads are added. Lightweight primitives such as EvalAdd show only modest changes across all counters. As homomorphically encrypted kernels continue to evolve, CryptOracle will offer detailed insights into their architectural challenges.
@@ 5.1.2 Microbenchmark Analysis
Figure 8 depicts the runtime scaling of matrix multiplication. When scaling up the ring dimension (from to ), the runtime grows almost linearly; doubling the ring dimension roughly doubles the runtime. Holding the ring dimension fixed and enlarging the matrix size also yields a linear rise, for instance an approximately 2.25 slowdown when the ring dimension is 17 with one thread (the highest point on the blue curve). This linear behavior under FHE contrasts with the quadratic scaling characteristic of plaintext matrix multiplication. Finally, as the thread count increases, larger matrices reach saturation sooner, causing the corresponding curves to converge.
@@ 5.2.1 Runtime and Energy Prediction
Figure 9 compares CryptOracle ’s predictions with measured results. There are two observations in Figure 9(a). First, in general, exhibits better prediction accuracy compared to workloads, since microbenchmarks have been well optimized for FHE challenges [24], while workloads are more for functional evaluations. Second, larger thread counts gives better accuracy, due to the performance saturates more. The geomean of prediction accuracy in dashed lines concurs with the two observations above. The overall geomean across all thread counts is -7.02% and 8.40% for microbenchmark and workload runtime, and -9.74% and 15.67% for microbenchmark and workload energy. We further show the correlation of our prediction in Figure 9(b). We observe consistently high cosine similarities, with chi-square giving the lowest correlation among all, owing to the large deviation between its primitive profiling results and actual runtime.
@@ 5.2.2 Prediction Breakdown
Figure 10 shows the contribution of primitives, expressed as a percentage of runtime or energy. Figure 10 (a) shows the predicted vs actual runtime breakdown for various workloads, while Figure 10 (b) contrasts predicted runtime vs predicted energy contributions. Each bar is stacked by the type of primitive kernels used in the application. Though we show a thorough breakdown, there is scope for "other" newer kernels to be supported by CryptOracle for future workloads. Figure 10 (a) indicates strong alignment between predicted and actual runtime across microbenchmarks. However, this deviates more for workloads, such as cifar10, resnet and logreg, due to their more complex nature. However, the prediction preserves the relative importance of primitives, i.e., operations that dominate the actual runtime also account for a proportionally larger share in the predicted runtime. Figure 10 (b) highlights the difference between predicted runtime and predicted energy, with most applications showing a similar distribution for runtime and energy.
Resnet exhibits a slight deviation, with EvalChebyshevFunction occupying a higher energy ratio than runtime. Thus, our results indicate that CryptOracle accurately preserves the relative contribution of individual primitives, capturing consistent trends in both runtime and energy profiles.
@@ 5.2.3 Prediction Overhead
We present the prediction overhead of CryptOracle in Table 8. The table compares the runtime analysis time, which measures execution time and energy without collecting performance counters, against the model prediction time, an almost constant cost incurred by running the Python-based performance model. Profiling latency varies considerably across workloads, and Intel CPUs consistently outperform AMD CPUs because of their larger on-chip memory capacities, in agreement with the primitive and microbenchmark results. The performance model accelerates runtime and energy analysis by , depending on the workload.
@@
5.2.4 Case Study
To illustrate another application of CryptOracle, we estimate the runtime and energy savings that can be achieved through algorithmic innovations. Specifically, we count the FHE primitive operations required by two algorithms implementing ciphertext–ciphertext matrix multiplication [50]. The projected gains are presented in Figure 11. Across various CPU vendors and thread counts, the projected speedups and energy reductions are consistent, ranging from roughly . The original study in [50] evaluated its implementation on a GPU; the measured speedup there is approximately , which differs from our CPU-based projection.
6 Limitation of This Work
We identify three primary limitations in this work, related to hardware support, benchmark optimization, and prediction methodology. First, CryptOracle currently supports only CPU platforms because of the limited scope of the OpenFHE backbone. Nevertheless, this limitation does not diminish the merit of CryptOracle; it remains a universally accessible, shared platform for examining advances in FHE applications and algorithms from both performance and energy perspectives. Second, this work does not explore any optimizations on the benchmark collections. Consequently, the predicted runtime and energy for full workloads exhibit greater variation than with the microbenchmarks, which have been used in FHE programming challenges [24]. Third, CryptOracle does not yet predict performance or efficiency on unseen hardware or algorithm configurations. This gap could be addressed by machine learning based techniques instead of simple linear extrapolation. Similar strategies have been extensively studied for AI workloads [48, 30, 39], and we leave their adoption to future work.
@@ 7 Related Work
7.1 FHE characterization
Although traditional benchmarks suites such as SPEC CPU2017 [11], MLPerf [42], and Graph500 [27] have been extensively studied, systematic performance evaluation of FHE remains sparse, but on the horizon.
Pal et al.[45] present a foundational study using OpenFHE[6], SEAL [53], and HEaaN [15], revealing runtime and energy trade offs across security levels, polynomial degrees, and threading models. FHEBench [31] supplies a standardized suite for latency and memory comparison across libraries, and Bergerat et al.[8] explore parameter optimization for TFHE.
HEBench [25], created by Intel and collaborators, evaluates canonical workloads such as dot products and matrix multiplication through a modular backend loader. Despite its robustness, it currently supports only a small number of libraries. HEProfiler [55] offers cycle accurate profiling of FHE primitives, exposing threading and memory inefficiencies. The native benchmark suite in OpenFHE [6] provides only basic profiling, limited to runtime, CPU core, and frequency statistics. It omits advanced metrics such as latency, power, and throughput, and it lacks profiling facilities such as FlameGraphs or runtime tracing that are crucial for comprehensive performance analysis.
@@ 7.1 FHE characterization
7.2 FHE acceleration
In recent years, accelerating FHE through both hardware and software has become a major research focus, driven by the wider adoption of privacy-preserving computation in machine learning and data analytics.
On the software side, frameworks such as Orion [21] and Cheetah [49] automatically optimize FHE programs by selecting efficient operator-packing strategies, managing rotation schedules, and reducing the need for bootstrapping. Cheetah reports speedups of up to 79 through compiler-level transformations, while Orion demonstrates practical viability by evaluating deep networks like ResNet-50.
At the same time, hardware research targets faster execution of the fundamental FHE primitives.
FPGA-based solutions such as FAB [3], FAST [56], and HEAP [2] exploit parallelism, on-chip data locality, and hybrid scheme switching to accelerate bootstrapping. ASIC-based designs like F1 [51], BTS [35], and CraterLake [52] integrate wide vector modular arithmetic units, optimized NTT hardware, and large on-chip memories. TensorFHE [23] exploits GPU Tensor Core Units to execute NTT and MULT operations efficiently in batches.
However, prior work concentrates mainly on reducing execution time and lacks systematic methodologies for profiling performance across workload types, security levels, and backend configurations. To address this gap, CryptOracle offers an analytical and empirical characterization framework for OpenFHE. It enables detailed analysis of runtime behavior, resource utilization, and the architectural impact of FHE applications, going beyond mere acceleration to deliver insights that are critical for co-design across software and hardware layers.
@@ 8 Conclusion
As privacy-preserving machine learning becomes increasingly critical, Despite the strong theoretical security guarantees, FHE has considerable computational overhead, restricting practical deployment. We introduced CryptOracle, a modular characterization framework that integrates a benchmark suite, a hardware profiler, and a performance model. This framework is built on top of OpenFHE for the CKKS scheme, enabling systematic evaluation across abstraction levels, hardware platforms and security configurations. By predicting runtime and energy consumption, CryptOracle highlights the performance trade-offs in FHE and fosters collaboration across application, algorithm, software, and hardware development. Our open-source release of CryptOracle aims to accelerate progress toward making FHE practical for real-world machine learning workloads.

### v3

@@ 4.1 Software Setup
To understand the trade-offs between computational efficiency and security, we sweep the ciphertext ring dimension N across and the computational depth L across [10, 20], maintaining a constant batch size B. We also profile the predefined security standards (128, 192, and 256-bit security from [HomomorphicEncryptionSecurityStandard]), which override the default parameter selection. To ensure a fair comparison, security parameters, evaluation keys, and input ciphertexts are generated once, serialized to files, and then reused by all primitives and microbenchmarks. Each workload fixes any additional application-specific parameters. Table 2 lists the default settings used in the evaluation, unless stated otherwise. During data collection, the median of five runs is reported for each metric. For primitive profiling, where an individual operation may complete in tens of microseconds, we invoke each primitive repeatedly until the cumulative runtime exceeds 500ms, then compute per-call averages based on the total number of invocations.
@@ 5.1.1 Primitive Analysis
Figure 6 explores the microarchitectural behaviour of the primitives on the AMD platform, presenting a subset of the events listed in Table 3. At low thread counts, EvalMult and EvalRotate record the largest instruction footprints, underscoring their computational intensity. As the thread count rises from one to four, the instruction counts drop substantially, most notably for EvalMult, indicating that multithreading removes redundant serial work. In contrast, cache references increase with additional threads because of greater data sharing. The pattern of dTLB loads follows the cache trend, as increased parallel memory accesses raise TLB pressure and introduce page-level contention. Correspondingly, page-fault rates fall as more threads are added. Lightweight primitives such as EvalAdd show only modest changes across all counters. As homomorphic encryption kernels continue to evolve, CryptOracle will offer detailed insights into their architectural challenges.
5.1.2 OpenFHE Hardware Backends
While this work explores profiling OpenFHE on CPUs, OpenFHE is designed as a modular library able to interface with GPUs and accelerators via a hardware backend, if available. Currently, the only officially-supported OpenFHE backend is the Homomorphic Encryption Accelerator Library (HEXL). This library accelerates integer arithmetic on devices supporting AVX512 instructions. Figure 8 compares an EvalMult operation’s runtime and energy consumption with and without HEXL acceleration, which both exhibit similar trends for increasing ring dimension. HEXL provides an average speedup of 1.8x, with a maximum speedup of 2.36x at , after which memory becomes the bottleneck in the operation at .
@@ 5.1.3 Microbenchmark Analysis
Figure 9 depicts the runtime scaling of matrix multiplication. When scaling up the ring dimension (from to ), the runtime grows almost linearly. Holding the ring dimension fixed and enlarging the matrix size also yields a linear rise, for instance an approximately 2.25 slowdown when the ring dimension is 17 with one thread (the highest point on the blue curve). This linear behavior under FHE contrasts with the quadratic scaling characteristic of plaintext matrix multiplication. Finally, as the thread count increases, larger matrices reach saturation sooner, causing the corresponding curves to converge.
@@ 5.2.1 Runtime and Energy Prediction
There are two observations in Figure 10. First, in general, the model exhibits better prediction accuracy compared to workloads, since microbenchmarks have been well optimized for FHE challenges [FhermaChallenges], while workloads are more for functional evaluations. Second, larger thread counts give better accuracy, as the performance saturates more. The geomean of prediction accuracy in dashed lines concurs with the two observations above. The overall geomean across all thread counts is -7.02% and 8.40% for microbenchmark and workload runtime, and -9.74% and 15.67% for microbenchmark and workload energy.
@@ 5.2.2 Prediction Breakdown
Figure 11 shows the contribution of primitives, expressed as a percentage of runtime or energy. Figure 11 (a) shows the predicted vs actual runtime breakdown for various workloads, while Figure 11 (b) contrasts predicted runtime vs predicted energy contributions. Each bar is stacked by the type of primitive kernels used in the application. Though we show a thorough breakdown, there is scope for "other" newer kernels to be supported by CryptOracle for future workloads. Figure 11 (a) indicates strong alignment between predicted and actual runtime across microbenchmarks. While this deviates more for workloads (e.g. cifar10, resnet and logreg) due to their more complex nature, the prediction preserves the relative importance of primitives. Figure 11 (b) highlights the difference between predicted runtime and predicted energy, with most applications showing a similar distribution for each. Resnet exhibits a slight deviation, with EvalChebyshevFunction occupying a higher energy ratio than runtime. Thus, our results indicate that CryptOracle accurately preserves the relative contribution of individual primitives, capturing consistent trends in both runtime and energy profiles.
@@ 5.2.3 Prediction Overhead
We present the prediction overhead of CryptOracle in Table 6. The table compares the runtime analysis time, which measures execution time and energy without collecting performance counters, against the model prediction time, which is an almost constant cost incurred by running the Python-based performance model. Profiling latency varies considerably across workloads, and Intel CPUs consistently outperform AMD CPUs because of their larger on-chip memory capacities, in agreement with the primitive and microbenchmark results. The performance model accelerates runtime and energy analysis by , depending on the workload.
@@
5.2.4 Case Study: Algorithmic Innovation
To illustrate another application of CryptOracle, we estimate the runtime and energy savings that can be achieved through algorithmic innovations. Specifically, we count the primitive operations required by two algorithms implementing ciphertext–ciphertext matrix multiplication [rho2024encryption]. The projected gains are presented in Figure 12. Across various CPU vendors and thread counts, the projected speedups and energy reductions are consistent, ranging from . The original study in [rho2024encryption] evaluated its implementation on a GPU; the measured speedup there is approximately , which differs from our CPU-based projection.
6 Limitations of This Work
We identify two primary limitations in this work. First, CryptOracle currently supports only CPU platforms due to the lack of supported OpenFHE hardware backends. This limitation does not diminish the merit of CryptOracle, as it remains an accessible platform for examining advances in FHE applications and algorithms. Also, GPU acceleration of FHE has shown great promise, and CryptOracle’s infrastructure can be expanded to support future GPU extensions to OpenFHE. Some of these extensions already exist [agullódomingo2025fideslibfullyfledgedopensourcefhe], but they are experimental and lack any full applications to benchmark. Second, CryptOracle does not yet predict performance or efficiency on unseen hardware or algorithm configurations. This gap could be addressed by machine learning-based techniques instead of simple linear extrapolation. Similar strategies have been extensively studied for AI workloads [qi17paleo, calculon_paper, 10.1145/3669940.3707265], and we leave their adoption to future work.
@@ 7 Related Work
Although traditional benchmarks suites such as SPEC CPU2017 [spec_cpu2017], MLPerf [mattson2020mlperftrainingbenchmark], and Graph500 [graph500] have been extensively studied, systematic performance evaluation of FHE remains sparse, but on the horizon.
Pal et al.[pal2023fully] present a foundational study using OpenFHE[OpenFHE], SEAL [sealcrypto], and HEaaN [cheon2017homomorphic], revealing runtime and energy trade-offs across security levels, polynomial degrees, and threading models. FHEBench [jiang2022fhebench] supplies a standardized suite for latency and memory comparison across libraries, and Bergerat et al.[bergerat2023parameter] explore parameter optimization for TFHE.
HEBench [hebench], created by Intel and collaborators, evaluates canonical workloads such as dot products and matrix multiplication through a modular backend loader. Despite its robustness, it currently supports only a small number of libraries. HEProfiler [takeshita2025heprofiler] offers cycle accurate profiling of FHE primitives, exposing threading and memory inefficiencies. The native benchmark suite in OpenFHE [OpenFHE] provides only basic profiling, limited to runtime, CPU core, and frequency statistics. It omits advanced performance metrics such as latency, power, and throughput and lacks event profiling, runtime tracing, and other profiling capabilities.
@@ 8 Conclusion
As privacy-preserving machine learning becomes increasingly critical, FHE offers strong theoretical security guarantees but also incurs considerable computational overhead, restricting practical deployment. We introduced CryptOracle, a modular characterization framework that integrates a benchmark suite, a hardware profiler, and a performance model to understand the runtime and energy consumption of CPUs running OpenFHE. Our open-source release of CryptOracle aims to accelerate progress toward making FHE practical for real-world machine learning workloads.